OpenStudy (anonymous):

served a ball at 155 miles per hour, the fastest serve ever recorded. What was the speed of the tennis ball, in feet per second, rounded to the nearest foot?

OpenStudy (anonymous):

Just a change of units. \[155\frac{miles}{hour} \times \frac{hour}{3600 \sec}\times \frac{5280 feet}{mile}=227\frac{1}{3}\frac{feet}{second }\approx227\frac{ft}{\sec}\]

OpenStudy (anonymous):

Note, every time you change the units, for example, by multiplying by\[\frac{hour}{3600\sec}\]we're really multiplying by 1 and therefore NOT changing the physical measurement. This is because, in this example, there are 3600 seconds in an hour, so we equate \[1h=3600s \rightarrow \frac{1h}{3600s}=1\]after dividing both sides by 2600 seconds. Hope it helps.